OpenStudy (anonymous):

I have the answer can someone please see if I did this right.... Suppose we have partially filled water tank and water begins entering the tank. The amount of water, in gallons in the tank is given by w(t), where t is the number of minutes after water began entering the tank. The graph above is of w'(t), the rate at which water is entering the tank in gallons/minute. Find and interpret

OpenStudy (anonymous):

After intrepret is this\[\int\limits_{0}^{20} w'(t) dt\]

OpenStudy (e.mccormick):

So what answer did you get?

OpenStudy (anonymous):

I got \[\int\limits_{0}^{20} 60d20\]

OpenStudy (anonymous):

then 0 as a final answer but I think I might have messed something up

OpenStudy (e.mccormick):

Well, because it is a integral of the derivative, the goal is to find the original equation.

OpenStudy (anonymous):

The original equation is the one that I posted first

OpenStudy (e.mccormick):

That derivative seems to be piecewise, which makes it a little more complex. It says the graph is of w'(t).

OpenStudy (anonymous):

right the left axis going vertical is that

OpenStudy (e.mccormick):

If you look, for \(0\le t < 6\) you have one rate, \(6\le t < 14\) is another, and \(14\le t < 20\) is a third equation because while it is similar to the first, the base is different.

OpenStudy (anonymous):

ok then what is next I guess I am confussed on how to solve the problem then

OpenStudy (anonymous):

I understand where you got the last post from what do I do next?

OpenStudy (e.mccormick):

Well, find a set of equations that would result in that graph. that will be w'(t). Then integrate it.

OpenStudy (anonymous):

how do you do that

OpenStudy (e.mccormick):

Which part of that is the issue?

OpenStudy (anonymous):

finding the equatino the next step

OpenStudy (e.mccormick):

OK. Well, those are linear representation, so we can use simple point slope. For \(0\le t < 6\) you can put any points in the form of \((t,w'(t))=(x,y)\). You have the points \((0,0)\) and \((2,10)\) that are on that line. So what would the equation of the line be?

OpenStudy (e.mccormick):

Yes, we are reviewing point/slope formula.... welcome to calculus.

OpenStudy (anonymous):

so then is the equation y=10x=0

OpenStudy (anonymous):

sorryy=10x+0

OpenStudy (anonymous):

@e.mccormick is that close at all I havenot done the point slope formula in a while

OpenStudy (anonymous):

are you still there

OpenStudy (e.mccormick):

Got called away for a bit.... I am at work. Hehe.

OpenStudy (anonymous):

oh ok sorry

OpenStudy (e.mccormick):

5x, not 10. 2t=10, so...

OpenStudy (anonymous):

y=5x+0

OpenStudy (e.mccormick):

So the first part of the piecewise function is: 5t for \(0\le t < 6\) The second part is really simple: 30 for \(6\le t < 14\) What would this leave for the last part? It is similar to the first part.

OpenStudy (anonymous):

70

OpenStudy (e.mccormick):

Well, it would still have the 5t in there somehow, that is why I say it is similar, but the starting t is 14 to get 30.... and 30=6*5, not 14*5.... sooooo....

OpenStudy (anonymous):

so then 16

OpenStudy (e.mccormick):

5(t-??)=something. That is the form it needs. And the first value needs to be: w'(14)=30

OpenStudy (anonymous):

ok would it be 2

OpenStudy (e.mccormick):

Lets try going back to point slope... you have these points: (14,30) and (16,40) Slope = \(\Delta y/\Delta x=(40-30)/(16-14)=5\) That much we know. Then, we take one point and find the formula using:\(y-y_1=m(x-x_1)\) where m=5, and \((x_1,y_1)\) is one of those two points.

OpenStudy (e.mccormick):

Yah, this is like Algebra 1 or 2, so by the time you are in calculus you have not seen these problems in a year or two. So it catches a lot of people off guard. Calculus is 90% doing algebra, trig, and so on to both set up and simplify the problems. \(y-30=5(x-14) \implies y-30=5x-70 \implies y=5x-40\) Remember doing those?

OpenStudy (e.mccormick):

This makes the piecewise function of:\[ w'(t) = \left\{ \begin{array}{lcl} 5t & \mathrm{for} &0\le t < 6\\ 30 & \mathrm{for} &6\le t < 14\\ 5t-40 & \mathrm{for} &14\le t \le 20 \end{array} \right. \]Now you need to integrate that. Remember that each piece will have its own integration rang based off of where it is on t. For example, instad of doing one integral from 0 to 20, the first part will be an integral of 0 to 6 for the first piece. THe second will be defined by the second piece, and the third by the third. I hope that gives you enough to finish whenever you get back.

OpenStudy (anonymous):

can someone please help me with this I have no idea what I am doing

OpenStudy (e.mccormick):

Ah ha! You were back...

OpenStudy (e.mccormick):

Now, you see the w'(t) U wrote there? That is what all the algebra was for. That is what you need to integrate o get w(t). But you can't do it in one shot.